OpenStudy (anonymous):

two angles are supplementary if their sum is180.the larger angle measure four degrees more then seven times the measure of a small angle.if x represents the measure of the smaller angle and these two angle are supplementary angle find the measure of each angle. the smaller measure the larger measure

OpenStudy (anonymous):

help me please

OpenStudy (ash2326):

We are given the smaller angle as x . Let the larger angle be y As x and y are supplementary so \[x+y=180.............................(1)\] the larger angle measure four degrees more then seven times the measure of a small angle.so \[y=7x+4....................(2)\]From first equation we get \[y=180-x\] Let's substitute this in second equation we get \[180-x=7x+4\] @ariana-help can you find x now?

OpenStudy (anonymous):

i dont get it

OpenStudy (ash2326):

@ariana-help tell me which part you don't get?

OpenStudy (anonymous):

wat do i got to muli the numbers with

OpenStudy (ash2326):

Did you understand how I got the last equation?

OpenStudy (anonymous):

wat the answer

Directrix (directrix):

Let x be the smaller angle and 180 - x the larger of the two supplementary angles. 180 - x = 7x + 4 180 - 4 = 7x + x 176 = 8x 22 = x ------> smaller measure 180 - x = 158 ----> larger measure

OpenStudy (anonymous):

it wrong

OpenStudy (paxpolaris):

22degrees & 158 degrees are the right answer.

OpenStudy (anonymous):

no it was not

OpenStudy (anonymous):

130 and 50

Directrix (directrix):

Post your work and answers, please. I would like to compare it to mine to pinpoint my error. Thanks.

Directrix (directrix):

the larger angle measure four degrees more then seven times the measure of a small angle. --> from the problem as posted Your answers -->130 and 50 Checking your work: the larger angle measure four degrees more then seven times the measure of a small 130 = 4+ 7(50) 130 = 4 + 350 130 = 354 Your answer is NOT correct. Check to see if the problem is correctly posted.